# Mathematical Language Models: A Survey

### v3

@@ 7. Conclusions
  • Hales et al. (2017) Thomas Hales, Mark Adams, Gertrud Bauer, Tat Dat Dang, John Harrison, Hoang Le Truong, Cezary Kaliszyk, Victor Magron, Sean McLaughlin, Tat Thang Nguyen, et al. 2017. A formal proof of the Kepler conjecture. In Forum of mathematics, Pi, Vol. 5. Cambridge University Press, e2.
  • Han et al. (2022) Jesse Michael Han, Jason Rute, Yuhuai Wu, Edward Ayers, and Stanislas Polu. 2022. Proof Artifact Co-Training for Theorem Proving with Language Models. In International Conference on Learning Representations.
  • Hao et al. (2023a) Shibo Hao, Yi Gu, Haodi Ma, Joshua Jiahua Hong, Zhen Wang, Daisy Zhe Wang, and Zhiting Hu. 2023a. Reasoning with language model is planning with world model. arXiv preprint arXiv:2305.14992 (2023).
  • Hao et al. (2023b) Shibo Hao, Tianyang Liu, Zhen Wang, and Zhiting Hu. 2023b. ToolkenGPT: Augmenting Frozen Language Models with Massive Tools via Tool Embeddings. arXiv:2305.11554 [cs.CL]
  • Hao et al. (2022) Yihan Hao, Mingliang Zhang, Fei Yin, and Linlin Huang. 2022. PGDP5K: A Diagram Parsing Dataset for Plane Geometry Problems. 2022 26th International Conference on Pattern Recognition (ICPR) (2022), 1763–1769.
  • He et al. (2022) Hangfeng He, Hongming Zhang, and Dan Roth. 2022. Rethinking with Retrieval: Faithful Large Language Model Inference. http://arxiv.org/abs/2301.00303 arXiv:2301.00303 [cs].
  • He-Yueya et al. (2023) Joy He-Yueya, Gabriel Poesia, Rose E Wang, and Noah D Goodman. 2023. Solving math word problems by combining language models with symbolic solvers. arXiv preprint arXiv:2304.09102 (2023).
  • Hendrycks et al. (2021a) Dan Hendrycks, Collin Burns, Steven Basart, Andy Zou, Mantas Mazeika, Dawn Song, and Jacob Steinhardt. 2021a. Measuring Massive Multitask Language Understanding. Proceedings of the International Conference on Learning Representations (ICLR) (2021).
  • Hendrycks et al. (2021b) Dan Hendrycks, Collin Burns, Saurav Kadavath, Akul Arora, Steven Basart, Eric Tang, Dawn Song, and Jacob Steinhardt. 2021b. Measuring Mathematical Problem Solving With the MATH Dataset. In Thirty-fifth Conference on Neural Information Processing Systems Datasets and Benchmarks Track (Round 2).
@@
  • Hosseini et al. (2014) Mohammad Javad Hosseini, Hannaneh Hajishirzi, Oren Etzioni, and Nate Kushman. 2014. Learning to Solve Arithmetic Word Problems with Verb Categorization. In Proceedings of the 2014 Conference on Empirical Methods in Natural Language Processing (EMNLP), Alessandro Moschitti, Bo Pang, and Walter Daelemans (Eds.). Association for Computational Linguistics, Doha, Qatar, 523–533. https://doi.org/10.3115/v1/D14-1058
  • Hsieh et al. (2023) Cheng-Yu Hsieh, Si-An Chen, Chun-Liang Li, Yasuhisa Fujii, Alexander Ratner, Chen-Yu Lee, Ranjay Krishna, and Tomas Pfister. 2023. Tool Documentation Enables Zero-Shot Tool-Usage with Large Language Models. arXiv:2308.00675 [cs.CL]
  • Huang et al. (2019) Daniel Huang, Prafulla Dhariwal, Dawn Song, and Ilya Sutskever. 2019. GamePad: A Learning Environment for Theorem Proving. In International Conference on Learning Representations.
  • Huang et al. (2016) Danqing Huang, Shuming Shi, Chin-Yew Lin, Jian Yin, and Wei-Ying Ma. 2016. How well do computers solve math word problems? large-scale dataset construction and evaluation. In Proceedings of the 54th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). 887–896.
  • Imani et al. (2023) Shima Imani, Liang Du, and Harsh Shrivastava. 2023. Mathprompter: Mathematical reasoning using large language models. arXiv preprint arXiv:2303.05398 (2023).
  • Irving et al. (2016) Geoffrey Irving, Christian Szegedy, Alexander A Alemi, Niklas Eén, François Chollet, and Josef Urban. 2016. Deepmath-deep sequence models for premise selection. Advances in neural information processing systems 29 (2016).
  • Jelassi et al. (2023) Samy Jelassi, Stéphane d’Ascoli, Carles Domingo-Enrich, Yuhuai Wu, Yuanzhi Li, and François Charton. 2023. Length Generalization in Arithmetic Transformers. https://doi.org/10.48550/arXiv.2306.15400
@@
  • Jiang et al. (2021) Albert Qiaochu Jiang, Wenda Li, Jesse Michael Han, and Yuhuai Wu. 2021. LISA: Language models of ISAbelle proofs. In 6th Conference on Artificial Intelligence and Theorem Proving. 378–392.
@@
  • Jiang et al. (2022) Albert Q. Jiang, Wenda Li, Szymon Tworkowski, Konrad Czechowski, Tomasz Odrzygóźdź, Piotr Miłoś, Yuhuai Wu, and Mateja Jamnik. 2022. Thor: Wielding Hammers to Integrate Language Models and Automated Theorem Provers. arXiv:2205.10893 [cs.AI]
@@
  • Jiang et al. (2023b) Albert Qiaochu Jiang, Sean Welleck, Jin Peng Zhou, Timothee Lacroix, Jiacheng Liu, Wenda Li, Mateja Jamnik, Guillaume Lample, and Yuhuai Wu. 2023b. Draft, Sketch, and Prove: Guiding Formal Theorem Provers with Informal Proofs. In The Eleventh International Conference on Learning Representations.
@@
  • Jiang et al. (2023a) Song Jiang, Zahra Shakeri, Aaron Chan, Maziar Sanjabi, Hamed Firooz, Yinglong Xia, Bugra Akyildiz, Yizhou Sun, Jinchao Li, Qifan Wang, et al. 2023a. Resprompt: Residual connection prompting advances multi-step reasoning in large language models. arXiv preprint arXiv:2310.04743 (2023).
@@
  • Jie et al. (2022) Zhanming Jie, Jierui Li, and Wei Lu. 2022. Learning to Reason Deductively: Math Word Problem Solving as Complex Relation Extraction. arXiv:2203.10316 [cs]
@@
  • Kaliszyk et al. (2017) Cezary Kaliszyk, François Chollet, and Christian Szegedy. 2017. HolStep: A Machine Learning Dataset for Higher-order Logic Theorem Proving. In International Conference on Learning Representations.
@@
  • Kalyan et al. (2021) Ashwin Kalyan, Abhinav Kumar, Arjun Chandrasekaran, Ashish Sabharwal, and Peter Clark. 2021. How Much Coffee Was Consumed During EMNLP 2019? Fermi Problems: A New Reasoning Challenge for AI. In Conference on Empirical Methods in Natural Language Processing (EMNLP 2021). Association for Computational Linguistics, 7318–7328.
@@
  • Kaplan et al. (2020) Jared Kaplan, Sam McCandlish, Tom Henighan, Tom B Brown, Benjamin Chess, Rewon Child, Scott Gray, Alec Radford, Jeffrey Wu, and Dario Amodei. 2020. Scaling laws for neural language models. arXiv preprint arXiv:2001.08361 (2020).
@@
  • Kenton and Toutanova (2019) Jacob Devlin Ming-Wei Chang Kenton and Lee Kristina Toutanova. 2019. BERT: Pre-training of Deep Bidirectional Transformers for Language Understanding. In Proceedings of NAACL-HLT. 4171–4186.
@@
  • Khalifa et al. (2023) Muhammad Khalifa, Lajanugen Logeswaran, Moontae Lee, Honglak Lee, and Lu Wang. 2023. Discriminator-Guided Multi-step Reasoning with Language Models. arXiv preprint arXiv:2305.14934 (2023).
@@
  • Kim et al. (2020) Bugeun Kim, Kyung Seo Ki, Donggeon Lee, and Gahgene Gweon. 2020. Point to the Expression: Solving Algebraic Word Problems using the Expression-Pointer Transformer Model. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP). Association for Computational Linguistics, Online, 3768–3779. https://doi.org/10.18653/v1/2020.emnlp-main.308
@@
  • Kim et al. (2022) Bugeun Kim, Kyung Seo Ki, Sangkyu Rhim, and Gahgene Gweon. 2022. EPT-X: An Expression-Pointer Transformer model that generates eXplanations for numbers. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). 4442–4458.
@@
  • Koncel-Kedziorski et al. (2015) Rik Koncel-Kedziorski, Hannaneh Hajishirzi, Ashish Sabharwal, Oren Etzioni, and Siena Dumas Ang. 2015. Parsing algebraic word problems into equations. Transactions of the Association for Computational Linguistics 3 (2015), 585–597.
  • Koncel-Kedziorski et al. (2016) Rik Koncel-Kedziorski, Subhro Roy, Aida Amini, Nate Kushman, and Hannaneh Hajishirzi. 2016. MAWPS: A math word problem repository. In Proceedings of the 2016 conference of the north american chapter of the association for computational linguistics: human language technologies. 1152–1157.
  • Kushman et al. (2014) Nate Kushman, Yoav Artzi, Luke Zettlemoyer, and Regina Barzilay. 2014. Learning to Automatically Solve Algebra Word Problems. In Proceedings of the 52nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Kristina Toutanova and Hua Wu (Eds.). Association for Computational Linguistics, Baltimore, Maryland, 271–281. https://doi.org/10.3115/v1/P14-1026
@@
  • Lample et al. (2022) Guillaume Lample, Timothee Lacroix, Marie-Anne Lachaux, Aurelien Rodriguez, Amaury Hayat, Thibaut Lavril, Gabriel Ebner, and Xavier Martinet. 2022. Hypertree proof search for neural theorem proving. Advances in Neural Information Processing Systems 35 (2022), 26337–26349.
@@
  • Lei et al. (2023) Bin Lei, Chunhua Liao, Caiwen Ding, et al. 2023. Boosting logical reasoning in large language models through a new framework: The graph of thought. arXiv preprint arXiv:2308.08614 (2023).
@@
  • Lewis et al. (2020) Mike Lewis, Yinhan Liu, Naman Goyal, Marjan Ghazvininejad, Abdelrahman Mohamed, Omer Levy, Veselin Stoyanov, and Luke Zettlemoyer. 2020. BART: Denoising Sequence-to-Sequence Pre-training for Natural Language Generation, Translation, and Comprehension. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics. 7871–7880.
@@
  • Lewkowycz et al. ([n. d.]) Aitor Lewkowycz, Anders Andreassen, David Dohan, Ethan Dyer, Henryk Michalewski, Vinay Ramasesh, Ambrose Slone, Cem Anil, Imanol Schlag, Theo Gutman-Solo, Yuhuai Wu, Behnam Neyshabur, Guy Gur-Ari, and Vedant Misra. [n. d.]. Solving Quantitative Reasoning Problems With Language Models. ([n. d.]).
@@
  • Li et al. (2023a) Guohao Li, Hasan Abed Al Kader Hammoud, Hani Itani, Dmitrii Khizbullin, and Bernard Ghanem. 2023a. CAMEL: Communicative Agents for” Mind” Exploration of Large Language Model Society. In Thirty-seventh Conference on Neural Information Processing Systems.
@@
  • Li et al. (2021) Wenda Li, Lei Yu, Yuhuai Wu, and Lawrence C Paulson. 2021. IsarStep: a Benchmark for High-level Mathematical Reasoning. In International Conference on Learning Representations.
@@
  • Li et al. (2023b) Yifei Li, Zeqi Lin, Shizhuo Zhang, Qiang Fu, Bei Chen, Jian-Guang Lou, and Weizhu Chen. 2023b. Making language models better reasoners with step-aware verifier. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). 5315–5333.
@@
  • Li et al. (2022) Zhongli Li, Wenxuan Zhang, Chao Yan, Qingyu Zhou, Chao Li, Hongzhi Liu, and Yunbo Cao. 2022. Seeking Patterns, Not Just Memorizing Procedures: Contrastive Learning for Solving Math Word Problems. arXiv:2110.08464 [cs]
@@
  • Liang et al. (2022) Zhenwen Liang, Jipeng Zhang, Lei Wang, Wei Qin, Yunshi Lan, Jie Shao, and Xiangliang Zhang. 2022. MWP-BERT: Numeracy-augmented Pre-Training for Math Word Problem Solving. arXiv:2107.13435 [cs]
@@
  • Lightman et al. (2023) Hunter Lightman, Vineet Kosaraju, Yura Burda, Harri Edwards, Bowen Baker, Teddy Lee, Jan Leike, John Schulman, Ilya Sutskever, and Karl Cobbe. 2023. Let’s Verify Step by Step. arXiv:2305.20050 [cs.LG]
@@
  • Ling et al. (2017) Wang Ling, Dani Yogatama, Chris Dyer, and Phil Blunsom. 2017. Program Induction by Rationale Generation: Learning to Solve and Explain Algebraic Word Problems. In Proceedings of the 55th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). 158–167.
@@
  • Ling et al. (2023) Zhan Ling, Yunhao Fang, Xuanlin Li, Zhiao Huang, Mingu Lee, Roland Memisevic, and Hao Su. 2023. Deductive Verification of Chain-of-Thought Reasoning. arXiv preprint arXiv:2306.03872 (2023).
@@
  • Liu et al. (2023a) Bo Liu, Yuqian Jiang, Xiaohan Zhang, Qiang Liu, Shiqi Zhang, Joydeep Biswas, and Peter Stone. 2023a. Llm+ p: Empowering large language models with optimal planning proficiency. arXiv preprint arXiv:2304.11477 (2023).
@@
  • Liu et al. (2022) Jiachang Liu, Dinghan Shen, Yizhe Zhang, Bill Dolan, Lawrence Carin, and Weizhu Chen. 2022. What Makes Good In-Context Examples for GPT-3?. In Deep Learning Inside Out: 3rd Workshop on Knowledge Extraction and Integration for Deep Learning Architectures, DeeLIO 2022. Association for Computational Linguistics (ACL), 100–114.
@@
  • Liu and Low (2023) Tiedong Liu and Bryan Kian Hsiang Low. 2023. Goat: Fine-tuned LLaMA Outperforms GPT-4 on Arithmetic Tasks. http://arxiv.org/abs/2305.14201 arXiv:2305.14201 [cs].
@@
  • Liu et al. (2019) Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov. 2019. RoBERTa: A Robustly Optimized BERT Pretraining Approach. arXiv:1907.11692 [cs]
@@
  • Liu et al. (2023b) Yixin Liu, Avi Singh, C. Daniel Freeman, John D. Co-Reyes, and Peter J. Liu. 2023b. Improving Large Language Model Fine-tuning for Solving Math Problems. http://arxiv.org/abs/2310.10047 arXiv:2310.10047 [cs].
@@
  • Long (2023) Jieyi Long. 2023. Large Language Model Guided Tree-of-Thought. arXiv preprint arXiv:2305.08291 (2023).
@@
  • Lu et al. (2021a) Pan Lu, Ran Gong, Shibiao Jiang, Liang Qiu, Siyuan Huang, Xiaodan Liang, and Song-chun Zhu. 2021a. Inter-GPS: Interpretable Geometry Problem Solving with Formal Language and Symbolic Reasoning. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers). 6774–6786.
@@
  • Lu et al. (2023a) Pan Lu, Baolin Peng, Hao Cheng, Michel Galley, Kai-Wei Chang, Ying Nian Wu, Song-Chun Zhu, and Jianfeng Gao. 2023a. Chameleon: Plug-and-Play Compositional Reasoning with Large Language Models. arXiv:2304.09842 [cs.CL]

### v4

@@ Acknowledgements.
  • Hales et al. (2017) Thomas Hales, Mark Adams, Gertrud Bauer, Tat Dat Dang, John Harrison, Hoang Le Truong, Cezary Kaliszyk, Victor Magron, Sean McLaughlin, Tat Thang Nguyen, et al. 2017. A formal proof of the Kepler conjecture. In Forum of mathematics, Pi, Vol. 5. e2.
  • Han et al. (2022) Jesse Michael Han, Jason Rute, Yuhuai Wu, Edward Ayers, and Stanislas Polu. 2022. Proof Artifact Co-Training for Theorem Proving with Language Models. In ICLR.
  • Han et al. (2024) Simeng Han, Hailey Schoelkopf, Yilun Zhao, Zhenting Qi, Martin Riddell, Wenfei Zhou, James Coady, David Peng, Yujie Qiao, Luke Benson, Lucy Sun, Alexander Wardle-Solano, Hannah Szabó, Ekaterina Zubova, Matthew Burtell, Jonathan Fan, Yixin Liu, Brian Wong, Malcolm Sailor, Ansong Ni, Linyong Nan, Jungo Kasai, Tao Yu, Rui Zhang, Alexander R. Fabbri, Wojciech Kryscinski, Semih Yavuz, Ye Liu, Xi Victoria Lin, Shafiq Joty, Yingbo Zhou, Caiming Xiong, Rex Ying, Arman Cohan, and Dragomir Radev. 2024. FOLIO: Natural Language Reasoning with First-Order Logic. In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing, EMNLP 2024, Miami, FL, USA, November 12-16, 2024, Yaser Al-Onaizan, Mohit Bansal, and Yun-Nung Chen (Eds.). Association for Computational Linguistics, 22017–22031. https://aclanthology.org/2024.emnlp-main.1229
  • Hao et al. (2023a) Shibo Hao, Yi Gu, Haodi Ma, Joshua Jiahua Hong, Zhen Wang, Daisy Zhe Wang, and Zhiting Hu. 2023a. Reasoning with language model is planning with world model. arXiv (2023).
  • Hao et al. (2023b) Shibo Hao, Tianyang Liu, Zhen Wang, and Zhiting Hu. 2023b. ToolkenGPT: Augmenting Frozen Language Models with Massive Tools via Tool Embeddings. arXiv:2305.11554
  • Hao et al. (2022) Yihan Hao, Mingliang Zhang, Fei Yin, and Linlin Huang. 2022. PGDP5K: A Diagram Parsing Dataset for Plane Geometry Problems. ICPR (2022), 1763–1769.
  • He et al. (2022) Hangfeng He, Hongming Zhang, and Dan Roth. 2022. Rethinking with retrieval: Faithful large language model inference. arXiv (2022).
  • He-Yueya et al. (2023) Joy He-Yueya, Gabriel Poesia, Rose E Wang, and Noah D Goodman. 2023. Solving math word problems by combining language models with symbolic solvers. arXiv (2023).
  • Hendrycks et al. (2021a) Dan Hendrycks, Collin Burns, Steven Basart, Andy Zou, Mantas Mazeika, Dawn Song, and Jacob Steinhardt. 2021a. Measuring Massive Multitask Language Understanding. ICLR (2021).
  • Hendrycks et al. (2021b) Dan Hendrycks, Collin Burns, Saurav Kadavath, Akul Arora, Steven Basart, Eric Tang, Dawn Song, and Jacob Steinhardt. 2021b. Measuring Mathematical Problem Solving With the MATH Dataset. In NeurIPS Datasets and Benchmarks Track (Round 2).
@@
  • Hosseini et al. (2024) Arian Hosseini, Xingdi Yuan, Nikolay Malkin, Aaron Courville, Alessandro Sordoni, and Rishabh Agarwal. 2024. V-star: Training verifiers for self-taught reasoners. arXiv (2024).
  • Hosseini et al. (2014) Mohammad Javad Hosseini, Hannaneh Hajishirzi, Oren Etzioni, and Nate Kushman. 2014. Learning to Solve Arithmetic Word Problems with Verb Categorization. In EMNLP. 523–533.
  • Hsieh et al. (2023) Cheng-Yu Hsieh, Si-An Chen, Chun-Liang Li, Yasuhisa Fujii, Alexander Ratner, Chen-Yu Lee, Ranjay Krishna, and Tomas Pfister. 2023. Tool Documentation Enables Zero-Shot Tool-Usage with Large Language Models. arXiv:2308.00675
  • Huang et al. (2019) Daniel Huang, Prafulla Dhariwal, Dawn Song, and Ilya Sutskever. 2019. GamePad: A Learning Environment for Theorem Proving. In ICLR.
  • Huang et al. (2016) Danqing Huang, Shuming Shi, Chin-Yew Lin, Jian Yin, and Wei-Ying Ma. 2016. How well do computers solve math word problems? large-scale dataset construction and evaluation. In ACL. 887–896.
  • Huang et al. (2024) Xijie Huang, Li Lyna Zhang, Kwang-Ting Cheng, Fan Yang, and Mao Yang. 2024. Fewer is More: Boosting Math Reasoning with Reinforced Context Pruning. In EMNLP. 13674–13695.
  • Hurst et al. (2024) Aaron Hurst, Adam Lerer, Adam P Goucher, Adam Perelman, Aditya Ramesh, Aidan Clark, AJ Ostrow, Akila Welihinda, Alan Hayes, Alec Radford, et al. 2024. Gpt-4o system card. arXiv (2024).
  • Imani et al. (2023) Shima Imani, Liang Du, and Harsh Shrivastava. 2023. Mathprompter: Mathematical reasoning using large language models. arXiv (2023).
  • Irving et al. (2016) Geoffrey Irving, Christian Szegedy, Alexander A Alemi, Niklas Eén, François Chollet, and Josef Urban. 2016. Deepmath-deep sequence models for premise selection. NeurIPS 29 (2016).
  • Jelassi et al. (2023) Samy Jelassi, Stéphane d’Ascoli, Carles Domingo-Enrich, Yuhuai Wu, Yuanzhi Li, and François Charton. 2023. Length generalization in arithmetic transformers. arXiv (2023).
@@
  • Jiang et al. (2021) Albert Qiaochu Jiang, Wenda Li, Jesse Michael Han, and Yuhuai Wu. 2021. LISA: Language models of ISAbelle proofs. In AITP. 378–392.
@@
  • Jiang et al. (2022) Albert Q. Jiang, Wenda Li, Szymon Tworkowski, Konrad Czechowski, Tomasz Odrzygóźdź, Piotr Miłoś, Yuhuai Wu, and Mateja Jamnik. 2022. Thor: Wielding Hammers to Integrate Language Models and Automated Theorem Provers. arXiv:2205.10893
@@
  • Jiang et al. (2023b) Albert Qiaochu Jiang, Sean Welleck, Jin Peng Zhou, Timothee Lacroix, Jiacheng Liu, Wenda Li, Mateja Jamnik, Guillaume Lample, and Yuhuai Wu. 2023b. Draft, Sketch, and Prove: Guiding Formal Theorem Provers with Informal Proofs. In ICLR.
  • Jiang et al. (2024) Dongwei Jiang, Marcio Fonseca, and Shay B Cohen. 2024. LeanReasoner: Boosting Complex Logical Reasoning with Lean. arXiv (2024).
@@
  • Jiang et al. (2023a) Song Jiang, Zahra Shakeri, Aaron Chan, Maziar Sanjabi, Hamed Firooz, Yinglong Xia, Bugra Akyildiz, Yizhou Sun, Jinchao Li, Qifan Wang, et al. 2023a. Resprompt: Residual connection prompting advances multi-step reasoning in large language models. arXiv (2023).
@@
  • Jie et al. (2022) Zhanming Jie, Jierui Li, and Wei Lu. 2022. Learning to Reason Deductively: Math Word Problem Solving as Complex Relation Extraction. arXiv:2203.10316
@@
  • Kaliszyk et al. (2017) Cezary Kaliszyk, François Chollet, and Christian Szegedy. 2017. HolStep: A Machine Learning Dataset for Higher-order Logic Theorem Proving. In ICLR.
@@
  • Kalyan et al. (2021) Ashwin Kalyan, Abhinav Kumar, Arjun Chandrasekaran, Ashish Sabharwal, and Peter Clark. 2021. How Much Coffee Was Consumed During EMNLP 2019? Fermi Problems: A New Reasoning Challenge for AI. In EMNLP. 7318–7328.
@@
  • Kaplan et al. (2020) Jared Kaplan, Sam McCandlish, Tom Henighan, Tom B Brown, Benjamin Chess, Rewon Child, Scott Gray, Alec Radford, Jeffrey Wu, and Dario Amodei. 2020. Scaling laws for neural language models. arXiv (2020).
  • Kazemi et al. (2022) Mehran Kazemi, Najoung Kim, Deepti Bhatia, Xin Xu, and Deepak Ramachandran. 2022. Lambada: Backward chaining for automated reasoning in natural language. arXiv (2022).
@@
  • Kenton and Toutanova (2019) Jacob Devlin Ming-Wei Chang Kenton and Lee Kristina Toutanova. 2019. BERT: Pre-training of Deep Bidirectional Transformers for Language Understanding. In NAACL. 4171–4186.
@@
  • Khalifa et al. (2023) Muhammad Khalifa, Lajanugen Logeswaran, Moontae Lee, Honglak Lee, and Lu Wang. 2023. Discriminator-Guided Multi-step Reasoning with Language Models. arXiv (2023).
@@
  • Kim et al. (2020) Bugeun Kim, Kyung Seo Ki, Donggeon Lee, and Gahgene Gweon. 2020. Point to the Expression: Solving Algebraic Word Problems using the Expression-Pointer Transformer Model. In EMNLP. Online, 3768–3779.
@@
  • Kim et al. (2022) Bugeun Kim, Kyung Seo Ki, Sangkyu Rhim, and Gahgene Gweon. 2022. EPT-X: An Expression-Pointer Transformer model that generates eXplanations for numbers. In ACL. 4442–4458.
@@
  • Koncel-Kedziorski et al. (2015) Rik Koncel-Kedziorski, Hannaneh Hajishirzi, Ashish Sabharwal, Oren Etzioni, and Siena Dumas Ang. 2015. Parsing algebraic word problems into equations. TACL 3 (2015), 585–597.
  • Koncel-Kedziorski et al. (2016) Rik Koncel-Kedziorski, Subhro Roy, Aida Amini, Nate Kushman, and Hannaneh Hajishirzi. 2016. MAWPS: A math word problem repository. In NAACL. 1152–1157.
  • Kumar et al. (2024) Aviral Kumar, Vincent Zhuang, Rishabh Agarwal, Yi Su, John D Co-Reyes, Avi Singh, Kate Baumli, Shariq Iqbal, Colton Bishop, Rebecca Roelofs, et al. 2024. Training language models to self-correct via reinforcement learning. arXiv (2024).
  • Kushman et al. (2014) Nate Kushman, Yoav Artzi, Luke Zettlemoyer, and Regina Barzilay. 2014. Learning to Automatically Solve Algebra Word Problems. In ACL. 271–281.
  • Lai et al. (2024) Xin Lai, Zhuotao Tian, Yukang Chen, Senqiao Yang, Xiangru Peng, and Jiaya Jia. 2024. Step-dpo: Step-wise preference optimization for long-chain reasoning of llms. arXiv (2024).
@@
  • Lample et al. (2022) Guillaume Lample, Timothee Lacroix, Marie-Anne Lachaux, Aurelien Rodriguez, Amaury Hayat, Thibaut Lavril, Gabriel Ebner, and Xavier Martinet. 2022. Hypertree proof search for neural theorem proving. NeurIPS 35 (2022), 26337–26349.
@@
  • Lei et al. (2023) Bin Lei, Chunhua Liao, Caiwen Ding, et al. 2023. Boosting logical reasoning in large language models through a new framework: The graph of thought. arXiv (2023).
@@
  • Lewis et al. (2020) Mike Lewis, Yinhan Liu, Naman Goyal, Marjan Ghazvininejad, Abdelrahman Mohamed, Omer Levy, Veselin Stoyanov, and Luke Zettlemoyer. 2020. BART: Denoising Sequence-to-Sequence Pre-training for Natural Language Generation, Translation, and Comprehension. In ACL. 7871–7880.
@@
  • Lewkowycz et al. (2022) Aitor Lewkowycz, Anders Andreassen, David Dohan, Ethan Dyer, Henryk Michalewski, Vinay Ramasesh, Ambrose Slone, Cem Anil, Imanol Schlag, Theo Gutman-Solo, et al. 2022. Solving quantitative reasoning problems with language models. NeurIPS 35 (2022), 3843–3857.
@@
  • Li et al. (2023a) Guohao Li, Hasan Abed Al Kader Hammoud, Hani Itani, Dmitrii Khizbullin, and Bernard Ghanem. 2023a. CAMEL: Communicative Agents for” Mind” Exploration of Large Language Model Society. In NeurIPS.
@@
  • Li et al. (2021) Wenda Li, Lei Yu, Yuhuai Wu, and Lawrence C Paulson. 2021. IsarStep: a Benchmark for High-level Mathematical Reasoning. In ICLR.
@@
  • Li et al. (2023b) Yifei Li, Zeqi Lin, Shizhuo Zhang, Qiang Fu, Bei Chen, Jian-Guang Lou, and Weizhu Chen. 2023b. Making language models better reasoners with step-aware verifier. In ACL. 5315–5333.
@@
  • Li et al. (2022) Zhongli Li, Wenxuan Zhang, Chao Yan, Qingyu Zhou, Chao Li, Hongzhi Liu, and Yunbo Cao. 2022. Seeking Patterns, Not Just Memorizing Procedures: Contrastive Learning for Solving Math Word Problems.
@@
  • Liang et al. (2022) Zhenwen Liang, Jipeng Zhang, Lei Wang, Wei Qin, Yunshi Lan, Jie Shao, and Xiangliang Zhang. 2022. MWP-BERT: Numeracy-augmented Pre-Training for Math Word Problem Solving.
@@
  • Lightman et al. (2024) Hunter Lightman, Vineet Kosaraju, Yuri Burda, Harrison Edwards, Bowen Baker, Teddy Lee, Jan Leike, John Schulman, Ilya Sutskever, and Karl Cobbe. 2024. Let’s Verify Step by Step. In ICLR.
@@
  • Ling et al. (2017) Wang Ling, Dani Yogatama, Chris Dyer, and Phil Blunsom. 2017. Program Induction by Rationale Generation: Learning to Solve and Explain Algebraic Word Problems. In ACL. 158–167.
@@
  • Ling et al. (2023) Zhan Ling, Yunhao Fang, Xuanlin Li, Zhiao Huang, Mingu Lee, Roland Memisevic, and Hao Su. 2023. Deductive Verification of Chain-of-Thought Reasoning. arXiv (2023).
@@
  • Liu et al. (2023a) Bo Liu, Yuqian Jiang, Xiaohan Zhang, Qiang Liu, Shiqi Zhang, Joydeep Biswas, and Peter Stone. 2023a. Llm+ p: Empowering large language models with optimal planning proficiency. arXiv (2023).
@@
  • Liu et al. (2022) Jiachang Liu, Dinghan Shen, Yizhe Zhang, Bill Dolan, Lawrence Carin, and Weizhu Chen. 2022. What Makes Good In-Context Examples for GPT-3?. In DeeLIO. 100–114.
@@
  • Liu and Low (2023) Tiedong Liu and Bryan Kian Hsiang Low. 2023. Goat: Fine-tuned llama outperforms gpt-4 on arithmetic tasks. arXiv (2023).
  • Liu et al. (2024) Wentao Liu, Qianjun Pan, Yi Zhang, Zhuo Liu, Ji Wu, Jie Zhou, Aimin Zhou, Qin Chen, Bo Jiang, and Liang He. 2024. Cmm-math: A chinese multimodal math dataset to evaluate and enhance the mathematics reasoning of large multimodal models. arXiv (2024).
@@
  • Liu et al. (2019) Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov. 2019. RoBERTa: A Robustly Optimized BERT Pretraining Approach.
@@
  • Liu et al. (2023b) Yixin Liu, Avi Singh, C Daniel Freeman, John D Co-Reyes, and Peter J Liu. 2023b. Improving large language model fine-tuning for solving math problems. arXiv (2023).
@@
  • Long (2023) Jieyi Long. 2023. Large Language Model Guided Tree-of-Thought. arXiv (2023).
  • Lu et al. (2023a) Pan Lu, Hritik Bansal, Tony Xia, Jiacheng Liu, Chunyuan Li, Hannaneh Hajishirzi, Hao Cheng, Kai-Wei Chang, Michel Galley, and Jianfeng Gao. 2023a. Mathvista: Evaluating mathematical reasoning of foundation models in visual contexts. arXiv (2023).
@@
  • Lu et al. (2021a) Pan Lu, Ran Gong, Shibiao Jiang, Liang Qiu, Siyuan Huang, Xiaodan Liang, and Song-chun Zhu. 2021a. Inter-GPS: Interpretable Geometry Problem Solving with Formal Language and Symbolic Reasoning. In ACL. 6774–6786.
@@
  • Lu et al. (2023b) Pan Lu, Baolin Peng, Hao Cheng, Michel Galley, Kai-Wei Chang, Ying Nian Wu, Song-Chun Zhu, and Jianfeng Gao. 2023b. Chameleon: Plug-and-Play Compositional Reasoning with Large Language Models. arXiv:2304.09842